Harry and Melissa Jacobson produce handcrafted furniture in a workshop on their farm. They have obtained

a load of 600 board feet of birch from a neighbor and are planning to produce round kitchen tables and ladder

back chairs during the next 3 months. Each table will require 30 hours of labor, each chair will require 18 hours,

and between them they have a total of 480 hours of labor available. A table requires 40 board feet of wood to

make, and a chair requires 15 board feet. A table earns the couple $575 in profit, and a chair earns $120 in profit.

Most people who buy a table also want four chairs to go with it, so for every table that is produced, at least four

chairs must also be made, although additional chairs can also be sold separately. Formulate and solve an integer

programming model to determine the number of tables and chairs the Jacobsons should make to maximize·profit.
Answer #1

ANSWER

First let us find out what is given to us :
Load acquired by Harry and Mellisa from neighbour board feet 600
Amount of Labor required by Table hours 30
Amount of Labor required by Chair hours 18
Total amount of labor available hours 480
Amount of board required by table board feet 40
Amount of board required by chair board feet 15
Profit on table USD 575
Profit on chair USD 120
Number of chairs produced for every table produced numbers 4
Now let the number of tables to be produced is represented by X1
& let the number of chairs to be produced are represented by X2
& let profit on entire business is represented by P
Now here the object function will be :
max P = X1*575+X2*120
Where constraints are :
X1*30+X2*18=480
X1*40+X2*15=600
X2=4*X1
X1 & X2 >0
Now let us use excel solver to solve this problem
Here we will test all available options and since number of chairs to be produced has to four times that of Tables produced our total calculation will be based upon the number of
tables produced . Now let us go ahead and produce 1 table and see the results.
X1 to be produced Numbers 1
X2 to be produced=4*X1 Numbers 4
Board required Board Consumed Labor required Labor Consumed Profit Total Profit
X A B=A*X C D=C*X p P*X
Board Feet/piece Board Feet Hours/piece Hours USD/Piece
X1 to be produced Numbers 1 40 40 30 30 575 23000
X2 to be produced Numbers 4 15 60 18 72 120 480
Total 100 102
Board & Labor hours available 600 480
Now in above we can see we have consumed only 100 board feet of board out of 680 and 102 hours of labor out of 480 , so let us try with 2 , 3 , 4 & 5 pieces of tables and stop where any of the constraints of board and labor hours is exceeded.
X1 to be produced Numbers 2
X2 to be produced=4*X1 Numbers 8
Board required Board Consumed Labor required Labor Consumed Profit Total Profit
X A B=A*X C D=C*X p P*X
Board Feet/piece Board Feet Hours/piece Hours USD/Piece
X1 to be produced Numbers 2 40 80 30 60 575 23000
X2 to be produced Numbers 8 15 120 18 144 120 960
Total 200 204
Board & Labor hours available 600 480
X1 to be produced Numbers 3
X2 to be produced=4*X1 Numbers 12
Board required Board Consumed Labor required Labor Consumed Profit Total Profit
X A B=A*X C D=C*X p P*X
Board Feet/piece Board Feet Hours/piece Hours USD/Piece
X1 to be produced Numbers 3 40 120 30 90 575 23000
X2 to be produced Numbers 12 15 180 18 216 120 1440
Total 300 306
Board & Labor hours available 600 480
X1 to be produced Numbers 4
X2 to be produced=4*X1 Numbers 16
Board required Board Consumed Labor required Labor Consumed Profit Total Profit
X A B=A*X C D=C*X p P*X
Board Feet/piece Board Feet Hours/piece Hours USD/Piece
X1 to be produced Numbers 4 40 160 30 120 575 23000
X2 to be produced Numbers 16 15 240 18 288 120 1920
Total 400 408 24920
Board & Labor hours available 600 480
X1 to be produced Numbers 5
X2 to be produced=4*X1 Numbers 20
Board required Board Consumed Labor required Labor Consumed Profit Total Profit
X A B=A*X C D=C*X p P*X
Board Feet/piece Board Feet Hours/piece Hours USD/Piece
X1 to be produced Numbers 5 40 200 30 150 575 23000
X2 to be produced Numbers 20 15 300 18 360 120 2400
Total 500 510 25400
Board & Labor hours available 600 480
Now as we can see we can produce a maximum of 4 tables and 16 chairs because if we are producing more than that then labor constraint is there.
Maximum profit we can earn is USD 24920

------------------------------------

DEAR STUDENT,

IF YOU HAVE ANY QUERY ASK ME IN THE COMMENT BOX,I AM HERE TO HELPS YOU.PLEASE GIVE ME POSITIVE RATINGS

*****************THANK YOU***************